OpenStudy (anonymous):

solve the system of equations: y = 2x x + 3y = 14

OpenStudy (anonymous):

nevermind i got it

OpenStudy (anonymous):

replace "y" in the second equation by "2x" and get one equation with one variable. you get \[x+3\times 2x=14\] \[x+6x=14\] and then solve for x

OpenStudy (anonymous):

x =2 y=4

OpenStudy (anonymous):

x=2 nd y=4